OpenStudy (anonymous):

simplify n^2+3n+2/2 divided by n+1/2(n+2)

OpenStudy (anonymous):

\[(n^{2}+3n+2)/2/(n+1)/2(n+2)=(n+1)(n+2)(2)(n+2)/2(n+1)=(n+2)^{2}\]